OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4).

OpenStudy (anonymous):

y = 4/3x + 8 y = 4/3x + 3 y = -3/4x + 8 y = -3/4x + 3

OpenStudy (alyssa_xo):

get it into point intercept form (y=mx+b) 3y=4x+12 y=\(\frac{4}{3}\)x+4 our line has the same slope, but passes through -3,4 we can use the point slope form equation for that. \(y-y_1=m(x-x_1)\)

OpenStudy (alyssa_xo):

do you need a little push?

OpenStudy (anonymous):

The correct answer is y=4/3 x +7

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I think u should check your option of question

OpenStudy (anonymous):

thats not an option

OpenStudy (alyssa_xo):

\(y_1\) is the y we want. \(x_1\) is the x we want \(y-y_1=m(x-x_1)\) m=\(\huge \frac{4}{3}\) \(y_1=4 | x_1=-3\)

OpenStudy (anonymous):

Then this is correct answer either there is mistake or there is some other point I means (-3,5) is this point in your book question check it.

OpenStudy (alyssa_xo):

so let's plug that stuff into the equation y-4=\(\frac{4}{3}\)(x--3) \(\huge y-4=\frac{4}{3}x-\cancel{3}*\frac{4}{\cancel{3}}\)

OpenStudy (alyssa_xo):

uh

OpenStudy (anonymous):

wait do i cross it off both sides

OpenStudy (alyssa_xo):

yeah, you sure it's (-3, 4).?

OpenStudy (anonymous):

yeah thats what it said im confused by it

OpenStudy (alyssa_xo):

no, not yet. that was just for \(\frac{4}{3}*3\)

OpenStudy (anonymous):

oh

OpenStudy (alyssa_xo):

god I'm so fucking stupid

OpenStudy (alyssa_xo):

\[\large y-4=\frac{4}{3}(x--3)\]\[\large y-4=\frac{4}{3}(x+3)\]

OpenStudy (anonymous):

do you think the question was written wrong or something

OpenStudy (alyssa_xo):

\[\large y-4=\frac{4}{3}x+\frac{4}{\cancel{3}}*\cancel{3})\]y-4=\(\frac{4}{3}x+4\)

OpenStudy (alyssa_xo):

y=4/3x+8

OpenStudy (alyssa_xo):

no, I'm just all over the place rn lol

OpenStudy (anonymous):

oh so y=4/3x+8 was the right answer????

OpenStudy (alyssa_xo):

yeah

OpenStudy (alyssa_xo):

do you see how we got there though?

OpenStudy (anonymous):

Your work helped me understnd these typeof problems alot better

OpenStudy (alyssa_xo):

noice, noice. to reiterate. step 1: get into point intercept form step 2: plug into point slope form